# Enhancing Action Recognition by Leveraging the Hierarchical Structure of Actions and Textual Context

### v1

@@ 3.2.1 Exploiting Hierarchical Information from Context
3.2.2 Fusion of Visual and Textual Embeddings
@@ 3.3 Training
4 Experiments
We now present the experimental setup on the Hierarchical TSU dataset. The results demonstrate the effectiveness of the proposed method compared to the equivalent visual-only approach, and state-of-the-art methods.
4.1 Hierarchical TSU Dataset
The experimental framework in this work is built upon the Hierarchical TSU dataset, a hierarchical dataset for activities of daily living, which extends the Toyota Smarthome Untrimmed (TSU) dataset [30]. The proposed dataset emphasizes the use of contextual information for action recognition, in contrast to the TSU dataset designed for action detection, and the original Toyota Smarthome dataset [61], which features short videos of individual actions with limited contextual information and a narrower number of classes.
The TSU dataset provides a rich variety of subjects and spatial settings within a home. With 51 different activity categories, it provides a comprehensive representation of daily living tasks. The dataset is captured by 7 static cameras at known locations, which provide additional contextual data. The combination of these cameras with the dense annotations provided for the untrimmed videos, makes this dataset a suitable option for our problem. In total, the dataset consists of 536 untrimmed videos with an average duration of 21 minutes. The videos capture the daily activities of 18 individuals aged between 60 and 80.
Among the 51 classes, we find elementary actions such as walk or get water, and composite ones that may overlap with the elementary ones, like make coffee or cook. Despite the hierarchical relationship between some composite and elementary classes, we maintain the original 51 classes from the TSU dataset and introduce a higher-level hierarchy comprising all action classes. We categorize these actions into 7 coarse-grained action categories based on the contextual meaning, location, and visual similarities, resulting in the Hierarchical TSU dataset. Each of the 51 categories is associated with one of the following classes: beverage preparation, general household activities, cleaning, prepare breakfast, use household appliances, cook and drink.
Figure 2: Example of annotations in the Hierachical TSU dataset. Video frames are labeled with fine-grained classes, such as composite activities (e.g., Make coffee). Each color represents a coarse-grained activity corresponding to its associated fine-grained class.
Action segments corresponding to action detection annotations are processed for contextual action recognition. From untrimmed annotations, each trimmed video corresponds to the duration of a fine-grained action segment, with possible overlaps both in fine- and coarse-grained actions due to composite activities, as depicted in Figure 2. Additionally, we process the sequence of previous actions to include longer temporal dependencies, which are treated as textual information. Finally, since the location of the fixed cameras is known and detailed in [30], we also include the location of the action as additional contextual information.
4.2 Experimental Setup
Evaluation: We evaluate the results of our experiments on top-$k$ accuracy following previous work on action recognition [7, 8, 61]. Top-$k$ accuracy measures how often the correct action label appears among the top-$k$ labels predicted by the model. We use the cross-subject evaluation introduced in the TSU dataset [30], which uses 11 subjects for training and the remaining 7 for testing. Unless specified otherwise, in all tables, the best performance is highlighted in bold, and the best performance within each subgroup, if applicable, is underlined.”
@@ 4.2 Experimental Setup
4.3 Results
To determine the best combination of visual input with contextual data that provides the best top-$k$ accuracy, Table 1 contains the results for the best settings obtained after extensive ablation experiments (see Section 5).
For each experiment, we evaluate using both a fine-grained only training objective (-), and a dual training objective with the proposed joint loss ($\checkmark$). The results lead to the conclusion that the proposed method outperforms the standard approach relying only on visual inputs. Moreover, for every input combination it is observed that predicting both fine- and coarse-grained actions results in better performance. Note that the observed decrease in top-5 accuracy in some cases is primarily due to the model’s high confidence in the top classes.
Table 1: Comparison of results: input modalities and joint loss on the best settings.
@@ 4.3 Results
Impact of optical flow on video analytics: Optical flow is considered an auxiliary modality that complements RGB data, providing a notable performance boost of $1.97\%$ when used in combination with RGB, and even outperforming in the uni-modal approach ($1.26\%$). Although, when combined with contextual information, its effectiveness diminishes, underperforming compared to the corresponding RGB-based method. When all three modalities (RGB, optical flow, and contextual information) are combined, optical flow contributes only a marginal improvement of $0.15\%$.
On the effect of the joint loss function: Using a joint loss function to predict both fine-grained and coarse-grained actions during training enhances performance across all input modalities. This approach yields up to a $1.37\%$ improvement when fusing RGB and text embeddings. Notably, the top-$k$ performance of the coarse-grained classifier is correlated with improvements in fine-grained action recognition.
To further validate the efficacy of the proposed method, we evaluate the best-performing model using contextual data extracted from its own predictions instead of ground truth annotations. Table 2 presents the top-$k$ accuracy for both fine- and coarse-grained actions. The results demonstrate the robustness of the method, even when relying on predicted labels. Compared to the previous joint loss results without contextual data, our method achieves a $2.08\%$ improvement with RGB data, a $2.03\%$ improvement with optical flow data, and a $2.92\%$ improvement with the combined RGB and optical flow approach.
Table 2: Action recognition performance when contextual data are obtained from predictions.
@@
4.4 Comparison with State-of-the-art Methods
To evaluate the reliability and robustness of our proposal, we compare our method with existing state-of-the-art architectures for action recognition. To this end, we train different models initialized to pre-trained Kinetics 400 [14] weights during $100$ epochs, the same number of epochs as the proposed method. 3D CNN models (VideoResNet [63], SlowFast [32] and X3D [64]) are trained with a learning rate of $10^{-4}$ and Adam optimizer, due to improved performance compared to the the training parameters of transformer based architectures. Such models (Video Swin [35] and MViT [37, 38]) are trained with the same hyperparameters as our model. Table 3 presents the results demonstrating that our model outperforms both 3D CNN and transformer based architectures on the Hierarchical TSU dataset, when trained with the same input size and hyperparameters. In this case the best performance, obtained from ground truth contextual data, is shown in bold. The second-best, derived from predictions, is underlined for each number of input blocks.
Table 3: Comparison to state-of-the-art methods trained on the Hierarchical TSU dataset.
@@ 4.4 Comparison with State-of-the-art Methods
5 Ablation Experiments
@@
We present an ablation study to evaluate the contributions of various components of the proposed method. In addition to identifying the optimal RGB feature extractor and determining the best configuration values for the transformers used in the architecture, we compare different methodologies for incorporating hierarchical information. We also conduct an ablation analysis on the influence of the number of past actions and location data within the contextual information. Furthermore, we experiment with GPT-3.5 and Llama 3 to investigate the effect of rephrasing on enriching contextual data, as well as the impact of different text encoders. Unless specified otherwise, for the experiments in this section we use the best configuration settings that yield a $54.95\%$ top-$1$ accuracy, and report results on fine-grained action recognition.
5.1 Contextual Data Analysis
@@
The best performing model considers the last 5 actions and the location provided by fixed cameras as contextual data. Text embeddings are obtained using DistilBERT [25] for feature extraction and the prompt used to generate the descriptions is a fixed template. Table 4 compares the use of DistilBERT and BERT as feature extractors as well as the use of rephrasing to enhance textual descriptions exploiting two well known large language models: GPT3.5 and Llama 3.
@@
Table 4: Effect of rephrasing and comparison of text feature extractors.
@@
Table 5: Study on the number of past actions and effect of location on contextual data.
@@
Results indicate that DistilBERT outperforms its predecessor for feature extraction. Additionally, a notable improvement is observed when using DistilBERT with a fixed template, as opposed to rephrasing for enhancing textual descriptions. Using the optimal configuration, we conducted a study to determine the best number of past actions and the impact of location on contextual descriptions. The results, presented in Table 5, suggest that five past actions is the optimal value on the Hierarchical TSU dataset. Regarding location information, we observe that it provides improvements only when using 1 or 5 past actions, with variability in other cases.
5.2 Strategies for Hierarchical Action Recognition
We compare various strategies to integrate fine-grained and coarse-grained action recognition into a joint learning scheme. Table 6 contains the results using 4 different mechanisms depending on the features used to disambiguate fine- and coarse-grained information. Specifically, (1) using only contextual information for coarse-grained action recognition. Another set of options is to use both visual and contextual information from the vision transformer, by (2) using two separate fusion transformers and (3) using two separate classifiers from the same fused embeddings. Another option (4) is to share partial layers of the classifier, separating only the last layers for classification.
Results indicate that the degree of differentiation between fine-grained and coarse-grained features significantly affects performance. Specifically, using only contextual information for coarse-grained action recognition (1) achieves the best performance, with a Top-1 accuracy of 54.95% for fine-grained actions and competitive results for coarse-grained actions. This confirms that contextual data alone is optimal for coarse-grained recognition and enhances fine-grained performance.
Table 6: Comparison of strategies for hierarchical action recognition.
@@ 5.2 Strategies for Hierarchical Action Recognition
  Contextual Data (1) 54.95 84.11 73.77 99.15
Separate Fusion (2) 51.83 82.14 73.10 99.19
Separate Classifier (3) 51.24 81.88 74.21 99.11
Shared Classifier (4) 48.17 78.55 73.18 99.44
5.3 Data Fusion
Effective fusion of data modalities is crucial for achieving strong model performance. To this end, we evaluate three different fusion strategies and conduct an ablation study on the most effective one: the fusion transformer. RGB and optical flow features are obtained separately, based on the finding that late fusion is the best strategy for the video encoder, as discussed in the next section. We assess the impact of three approaches: (1) concatenating all three modalities (RGB, optical flow, and text), (2) concatenating the visual modalities (RGB and optical flow) prior to the fusion transformer, and (3) inputting the three modalities separately. Table 7 contains the results, with visual concatenation prior to the fusion transformer performing best on top-1 accuracy.
Table 7: Comparison of fusion strategies.
@@ 5.3 Data Fusion
  Separate Modalities 49.94 80.03
Visual Concat 54.95 84.11
Concat 40.53 74.25
Since the fusion transformer outperforms the concatenation approach, we conduct an ablation study on the number of attention heads and encoder layers, while keeping the embedding size fixed at 768, which corresponds to the text modality. Table 9 presents the results for different numbers of encoder layers, with two layers yielding the best performance. Similarly, Table 9 shows the results for varying numbers of attention heads, with two heads being the optimal configuration for our method.
Table 8: Ablation study on the number of encoder layers on the fusion transformer.

### v2

@@ IV-B1 Exploiting Hierarchical Information from Context
IV-B2 Fusion of Visual and Textual Embeddings
@@ IV-C Training
V Experiments
This section presents the experimental setup and results to demonstrate the effectiveness of the proposed method compared to the equivalent visual-only approach, and state-of-the-art methods across different datasets.
V-A Experimental Setup
Evaluation: We evaluate the results of our experiments on top-$k$ accuracy following previous work on action recognition [10, 11, 67, 8, 34]. Top-$k$ accuracy measures how often the correct action label appears among the top-$k$ labels predicted by the model. We use the cross-subject evaluation introduced in the TSU dataset [33], which uses 11 subjects for training and the remaining 7 for testing. Unless specified otherwise, in all tables, the best performance is highlighted in bold, and the best performance within each subgroup, if applicable, is underlined.
@@ V-A Experimental Setup
V-B Results
To determine the best combination of visual input with contextual data that provides the best top-$k$ accuracy, Table II presents the results for the best settings obtained after extensive ablation experiments (see Section VI).
For each experiment, we evaluate using both a fine-grained-only training objective (-), and a dual training objective with the proposed joint loss ($\checkmark$). The results lead to the conclusion that the proposed method outperforms the standard approach relying only on visual inputs. Moreover, for every input combination it is observed that predicting both fine- and coarse-grained actions results in better performance. Note that the observed decrease in top-$5$ accuracy in some cases is primarily due to the model’s high confidence in the top classes.
Impact of optical flow on video analytics: Optical flow is considered an auxiliary modality that complements RGB data, providing a notable performance boost of $1.97\%$ when used in combination with RGB, and even outperforming in the uni-modal approach ($1.26\%$). However, its effectiveness diminishes when combined with contextual information and underperforms compared to the corresponding RGB-based method. When all three modalities (RGB, optical flow, and contextual information) are combined, optical flow contributes only a marginal improvement of $0.15\%$.
On the effect of the joint loss function: Using a joint loss function to predict both fine-grained and coarse-grained actions during training enhances performance across all input modalities. This approach yields up to a $1.37\%$ improvement when fusing RGB and text embeddings. Notably, the top-$k$ performance of the coarse-grained classifier correlates with improvements in fine-grained action recognition.
To further validate the efficacy of the proposed method, we evaluate the best-performing model using contextual data extracted from its own predictions instead of ground truth annotations. Table III presents the top-$k$ accuracy for both fine- and coarse-grained actions. The results demonstrate the robustness of the method, even when relying on inference results. Compared to the previous joint loss results without contextual data, our method achieves a $2.08\%$ improvement with RGB data, a $2.03\%$ improvement with optical flow data, and a $2.92\%$ improvement with the combined RGB and optical flow approach.
TABLE II: Comparison of results: input modalities and joint loss on the best settings.
@@ V-B Results
TABLE III: Action recognition performance when contextual data are obtained from predictions.
@@
V-C Comparison with State-of-the-art Methods
To evaluate the reliability and robustness of our approach, we compare our method with existing state-of-the-art architectures for action recognition. Given that this work introduces the Hierarchical TSU dataset, the results serve as a baseline to evaluate where our method stands relative to existing approaches. To this end, we train state-of-the-art methods for action recognition initialized with pre-trained Kinetics 400 [17] weights for $100$ epochs, matching the training duration of our proposed method. 3D CNN models, including VideoResNet [72], SlowFast [36], and X3D [73], are trained using the Adam optimizer with a learning rate of $10^{-4}$, as this configuration showed improved performance for these architectures. Transformer-based models, consisting of Video Swin [39], MViT [41, 42], TimeSformer [38] and ViVit [37], are trained using the same hyperparameters as our method for consistency. To provide a fair comparison across architectures, we evaluate performance using input sizes that match those required by the respective state-of-the-art models.
Table IV summarizes the findings, demonstrating that our model outperforms both 3D CNN and transformer-based architectures on the Hierarchical TSU dataset when trained under identical input size and hyperparameter settings. The best performance, achieved using ground truth contextual data, is highlighted in bold, while the second-best results, derived from inference contextual data, are underlined for each number of input blocks.
TABLE IV: Comparison to state-of-the-art methods trained on the Hierarchical TSU dataset.
@@ V-C Comparison with State-of-the-art Methods
V-D Evaluation across Datasets
We further evaluate the proposed method on two additional action recognition benchmarks, IkeaASM and Assembly101. These industrial-like datasets are chosen for their inclusion of key elements relevant to our approach: untrimmed videos that support contextual modeling of prior actions, and hierarchical action structures.
IkeaASM [34]: This dataset focuses on furniture assembly tasks in different settings comprising 33 atomic (fine-grained) actions under 4 high-level (coarse-grained) categories, each consisting of a different furniture type. Unlike the Hierarchical TSU, a single fine-grained class is assigned to multiple coarse-grained classes. We follow the default train/test data split using the top camera view to ensure comparability with prior work. The model architecture and feature extractors match the best settings on the Hierarchical TSU. Given the absence of known locations, we exclude location information from the prompt, and adapt the domain information to “a person building furniture”. Results are presented in Table V, following [34] we report Top-1/3 accuracy.
Assembly101 [8]: This large-scale dataset captures the procedural assembly and disassembly of toy vehicles. It includes 1380 fine-grained and 199 coarse-grained action classes. Similar to IkeaASM, the dataset allows a fine-grained action to be associated with multiple coarse-grained categories, as fine- and coarse-grained recognition tasks are treated independently. To establish fine-to-coarse relationships in this dataset, we matched corresponding temporal segments.
We use the default train/test split and employ the provided DINOv2 [74] RGB features from the v4 camera for comparison, which yields the best performance in the third-person view. Because the features are pre-extracted, optical flow is not used, otherwise, the model setup remains the same. As with IkeaASM, prompts omit location information, and the domain context is described as “a person assembling a toy.” The results are reported in Table VI.
TABLE V: Ablation and comparison to state-of-the-art methods on IkeaASM dataset. SOTA extracted from [34]. Visual: RGB + Flow. Contextual: RGB + Flow + Text.
  Model Hier? Top-1 Top-3
  Ours (Visual) - 75.77 83.68
75.43 84.54
Ours (Contextual) - 76.98 88.49
79.04 76.98
C3D [75] - 45.73 69.56
P3D [76] - 60.40 81.07
I3D [17] - 57.58 76.72
Multimodal (RGB, Pose, Depth) [34] - 64.02 81.45
Ours (Contextual from predictions) 65.29 81.96
TABLE VI: Ablation and comparison to state-of-the-art methods on Assembly101 dataset. SOTA extracted from [8]. Visual: RGB. Contextual: RGB + Text.
  Model Hier? Top-1 Top-5
  Ours (Visual) - 72.98 84.25
73.06 83.96
Ours (Contextual) - 81.44 89.80
81.89 89.41
TSM [77] - 38.3 -
TempAgg [78] - 40.5 -
Ours (Contextual from predictions) 43.47 69.48
Results on both benchmarks demonstrate the effectiveness of the proposed method over prior work (see the second half of Table V and Table VI). On IkeaASM, our method using contextual information from predictions achieves an improvement of nearly 5% in Top-1 accuracy over the previous state-of-the-art RGB-based method, and over 1% compared to a multimodal approach that combines RGB, depth and 3D pose data. On Assembly101, it yields almost a 3% gain in Top-1 accuracy compared to TempAgg [78]. The ablation studies in the first half of the same tables further highlight the impact of contextual modeling, showcasing the benefits of leveraging longer temporal cues and introducing domain-specific textual information.
In contrast, incorporating hierarchical structures results in only marginal improvement on Assembly101 and even a slight decrease in performance on IkeaASM. We attribute this drop to label ambiguity. In IkeaASM, fine-grained actions are commonly assigned to multiple, or even all, coarse-grained categories, making the hierarchy noisy and less informative. As previously discussed, the hierarchical structure in IkeaASM was not designed explicitly for hierarchical action recognition but rather to organize the data by furniture type. Similarly, in Assembly101, hierarchical labels are treated as a separate task rather than an intrinsic semantic structure. Consequently, instead of guiding the model, these hierarchies can introduce noise and hinder fine-grained recognition. This effect is more pronounced in IkeaASM and only mildly beneficial in Assembly101, where over 50% of coarse-grained classes share at least one fine-grained action, and only 25% of fine-grained actions are uniquely assigned. These results underscore the importance of a well-structured and semantically meaningful label hierarchy in action understanding datasets, as unclear or overlapping hierarchies can negatively impact performance.
V-E Computational Analysis
We conduct a comparative analysis of the computational characteristics with respect to the previously introduced baseline methods in Table VII. Compared to prior works, our model shows a larger learning capacity, reflected in a higher number of trainable parameters, while maintaining a lower inference cost. Methods based on CNN architectures have lighter computational requirements, but they generally under-perform compared to transformer-based approaches. In contrast, our method achieves substantially better accuracy than other transformer-based models while requiring less computations during inference. Specifically, our model comprises 143.06 million trainable parameters, augmented by an additional 86.39 million parameters from RGB feature extraction, 10.3 million from optical flow feature extraction and 66.36 million from the transformer textual encoder. These parameters were obtained using the feature extractors that performed best.
In terms of inference complexity, our approach requires 13.78 GFLOPs for the multimodal fusion and classification. The visual feature extraction cost is calculated as $(33.74+2.85)\times 32$ GFLOPs, accounting for both the RGB and optical flow pathways across 32 frames. Additionally, the textual encoding incurs a cost of 11.18 GFLOPs. Overall, this leads to a substantially lower computational cost compared to transformer-based methods like TimeSformer and ViVit, or CNN-based approaches such as SlowFast, which process a larger number of frames or use additional frames to capture richer temporal cues despite similar input window sizes.
TABLE VII: Computational requiremments of state-of-the-art methods (S: Small, B: Base).
  Method # Frames
Trainable
Params (M)
Total
Params (M)
GFlops
  X3D 16 3.08 3.08 10.04
MViT (B) 36.34 36.34 141.8
MViTv2 (S) 34.27 34.27 128.64
VideoResNet 32 33.19 33.19 25.44
SlowFast 34.6 34.6 1970
VideoSwin (S) 49.55 49.55 244.66
TimeSformer 121.4 121.4 1519.3
ViVit 88.69 88.69 906.52
Ours 143.06 306.11 610.4
VI Ablation Experiments
@@
We present an ablation study to assess the contributions of the individual components within our proposed method. In addition to identifying the optimal RGB feature extractor and determining the best configuration values for the transformers used in the architecture, we compare various approaches for incorporating hierarchical information. We also analyze the influence of the number of past actions and the inclusion of location data within the contextual information. Furthermore, we investigate the impact of rephrasing with GPT-3.5 and Llama 3 on enriching contextual data, as well as the effect of different text encoders. Unless specified otherwise, all experiments in this section are conducted using the best configuration settings, which result in a $54.95\%$ top-11 accuracy, and we report the outcomes on fine-grained action recognition.
VI-A Contextual Data Analysis
@@
The best performing model considers the last 5 actions and the location provided by fixed cameras as contextual data. Text embeddings are obtained using DistilBERT [28] for feature extraction and the prompt used to generate the descriptions is a fixed template. Table VIII compares the use of DistilBERT and BERT as feature extractors as well as the use of rephrasing to enhance textual descriptions exploiting two well known large language models: GPT3.5 and Llama 3.
@@
TABLE VIII: Effect of rephrasing and comparison of text feature extractors.
@@
TABLE IX: Study on the number of past actions and effect of location on contextual data.
@@
Results indicate that DistilBERT outperforms its predecessor for feature extraction. Additionally, a notable improvement is observed when using DistilBERT with a fixed template, as opposed to rephrasing for enhancing textual descriptions.
Based on the optimal configuration, we conducted a study to determine the best number of past actions and assess the impact of location on contextual descriptions. The results, presented in Table IX, suggest that using five past actions yields the best performance on the Hierarchical TSU dataset. Regarding location information, we observe that it provides improvements only when using 1 or 5 past actions, with variability in other cases.
VI-B Strategies for Hierarchical Action Recognition
As discussed in Section IV, fine-grained action recognition is performed using fused embeddings that combine both visual and contextual information. In addition, we compare different strategies for integrating fine-grained and coarse-grained action recognition into a joint learning framework. Table X presents the results of four mechanisms, each depending on the features used to disambiguate fine- and coarse-grained information. Specifically, the strategies are: (1) using only contextual information for coarse-grained action recognition; (2) using two separate fusion transformers, employing the same visual and contextual features as the fine-grained classifier; (3) using two separate classifiers from the same fused embeddings for fine- and coarse-grained classification; and (4) sharing part of the classifier layers, while separating only the final classification layers.
TABLE X: Comparison of strategies for hierarchical action recognition.
@@ VI-B Strategies for Hierarchical Action Recognition
  (1) Contextual Data 54.95 84.11 73.77 99.15
(2) Separate Fusion 51.83 82.14 73.10 99.19
(3) Separate Classifier 51.24 81.88 74.21 99.11
(4) Shared Classifier 48.17 78.55 73.18 99.44
TABLE XI: Comparison of fusion strategies.
@@
  (1) Separate Modalities 49.94 80.03
(2) Visual Concat 54.95 84.11
(3) Concat 40.53 74.25
The results indicate that the differentiation between fine-grained and coarse-grained features plays a significant role in performance. Specifically, using only contextual information for coarse-grained action recognition (1) achieves the best performance, with a top-1 accuracy of 54.95% for fine-grained actions and competitive results for coarse-grained actions. This suggests that contextual data is particularly effective for coarse-grained recognition and also helps to improve fine-grained action accuracy.
Although the use of contextual data improves coarse-grained recognition, it does not achieve the highest top-1 accuracy. However, it remains competitive compared to other approaches. We attribute this to the role of coarse-grained actions as an auxiliary modality that primarily enhances fine-grained performance, rather than being the main focus of the model. The highest top-1 accuracy for coarse-grained actions is achieved when incorporating visual information, as seen in the separate classifier strategy (3), which yields the best performance for coarse-grained recognition.
Furthermore, strategies (2)-(4), which involve increasing the number of shared layers between fine-grained and coarse-grained actions, lead to a reduction in fine-grained accuracy. We argue that this occurs because fine-grained actions rely on more detailed, discriminative visual features, while coarse-grained actions benefit from broader, higher-level visual cues.
VI-C Data Fusion
Effective fusion of data modalities is crucial for achieving strong model performance. To this end, we evaluate three different fusion strategies and conduct an ablation study on the most effective one: the fusion transformer. RGB and optical flow features are obtained separately, based on the finding that late fusion is the best strategy for the video encoder, as discussed in the next section. We assess the impact of three approaches: (1) concatenating all three modalities (RGB, optical flow, and text), (2) concatenating the visual modalities (RGB and optical flow) prior to the fusion transformer, and (3) inputting the three modalities separately. Table XI contains the results, with visual concatenation prior to the fusion transformer performing best on top-$1$ accuracy.
Since the fusion transformer outperforms the concatenation approach, we conduct an ablation study on the number of attention heads and encoder layers, while keeping the embedding size fixed at 768, which corresponds to the text modality. Table XII presents the results for different numbers of encoder layers, with two layers yielding the best performance. Similarly, Table XIII shows the results for varying numbers of attention heads, with two heads being the optimal configuration for our method.
TABLE XII: Ablation study on the number of encoder layers on the fusion transformer.